OpenStudy (anonymous):

What is the slope of the line that passes through (1, 4) and (1, −3)? 7 −7 undefined 0

OpenStudy (jack1):

(1, 4) and (1, −3) (x1,y1) and (x2, y2) m = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

m=(4-(-3))/(1-1) m=7/0 m= undefined

OpenStudy (jack1):

perfect, its a vertical line

OpenStudy (anonymous):

thanks @Jack1 AND @Aditya9826042721 help with another question?

OpenStudy (anonymous):

ya that s ok

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)? y − 1 = −2(x − 4) y − 1 = 2(x + 2) y − 13 = −2(x + 2) y − 13 = 2(x + 4)

OpenStudy (mathmale):

Please post each new question separately. Are you familiar with the point-slope formula for the equation of a straight line? You'll need this to answer this question.

OpenStudy (anonymous):

is not that que. u asked later?

OpenStudy (anonymous):

is it y=mx+b?

OpenStudy (mathmale):

that's the slope-intercept form. Please look up "point-slope form".

OpenStudy (anonymous):

not but (y-y1)=((y1-y2)/(x1-x2))(x-x1)

OpenStudy (mathmale):

OK, alesha: you are given two points on the line in question. First, find the slope of this line (see Aditya's formula) Second, insert either point, along with the slope, into the point slope formula. Simplify the result. That's your answer.

OpenStudy (anonymous):

ya we can just wait

OpenStudy (mathmale):

What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)? 1) Find the slope of this line. 2) Substitute the result (slope) and the coordinates of either given point into the point-slope form of the equaion of a straight line:\[y-y _{1}=m(x-x _{1})\]

OpenStudy (mathmale):

@aleisha96 : Please get involved: Calculate the slope on your own and display your work here.

OpenStudy (anonymous):



OpenStudy (mathmale):

@Aditya9826042721 : Thank you for all your inputs. Would you please GUIDE Aleshia through this problem, so that she can find the desired answers on her own? Please do not do the work for others; guide them in doing that work for themselves.

OpenStudy (anonymous):

ya u are right but today is my first day

OpenStudy (mathmale):